OpenStudy (anonymous):

The length of a rectangle is 6 meters more than its width. If the area is 135 square meters, what is its width? Please explain!

OpenStudy (anonymous):

area is length times width \(l\times w\) and \(l=w+6\) so it is \[A=(w+6)w\]

OpenStudy (anonymous):

you know that this is \(135\) solve \[w(w+6)=135\] for \(w\)

OpenStudy (anonymous):

can you solve that one?

OpenStudy (anonymous):

Yes. Thank you!!!

OpenStudy (anonymous):

yw